OpenStudy (anonymous):

Please help! Simplify the expression sqrt(250x) into the form Asqrt(B)

OpenStudy (anonymous):

\[250=25\times 10\] so \[\sqrt{250}=\sqrt{25\times 10}=\sqrt{25}\sqrt{10}=5\sqrt{10}\]

OpenStudy (anonymous):

Thanks! @satellite73